The 4th week of no wet mopping of the floor and no wet wiping the vacuum envelope.
We cleaned the intakes of the south arm IFO air condition. The bottom duct have quite a bit accumulation Atm1
See wet wiped contrast on Atm2
We found 3 holes around pipes ( coming from CES ) on the east wall that has to be sealed.
After closer examination of these holes, they are sealed off well.
Air condition maintenance will be finished by 11:30am tomorrow.
Yesterday, Kelvin mopped with chemicals the whole floor area of the lab. This was triggered by some visiting ants at our PSL last week. It was 6 months since we had the last fully wet mopped IFO floor.
The cleaning- mopping water became very dirty at the end.